OpenStudy (anonymous):

I NEED HELP PL CICK ON THE DOC1 im so lost

OpenStudy (anonymous):

IM LOST ON IT PL HELP

OpenStudy (anonymous):

PL I NEED HELP ON IT IM SO LOST

OpenStudy (anonymous):

I KNOW A is right but b and c im lost on

OpenStudy (anonymous):

pl help me out

OpenStudy (anonymous):

im so lost dam it

OpenStudy (wolf1728):

45 49 53 There are 4 more seats for each row so it is an arithmetic sequence seats in row1 = 41 + (row number * 4) row2 = 41 + (2 * 4) = 49 row3 = 41 + (3*4) = 53 Row 19 would be row 19 = 41 + (19*4) = 117

OpenStudy (anonymous):

got is b 19 = 41 + (19*4)

OpenStudy (anonymous):

in is one 117 is c

OpenStudy (anonymous):

im right

OpenStudy (wolf1728):

the answer to c is 117

OpenStudy (anonymous):

ok b is what

OpenStudy (wolf1728):

I'm not sure. They want the formula for the sequence row 'n' = 41 + (n * 4)

OpenStudy (anonymous):

ok it look like an=41+(n*4)

OpenStudy (wolf1728):

So, I guess you have it!

OpenStudy (anonymous):

ok thank you

OpenStudy (wolf1728):

c ya